# Fusing Bluetooth with Pedestrian Dead Reckoning: A Floor Plan-Assisted Positioning Approach

### v2

@@ Abstract
Floor plans can provide valuable prior information that helps enhance the accuracy of indoor positioning systems. However, existing research typically faces challenges in efficiently leveraging floor plan information and applying it to complex indoor layouts. To fully exploit information from floor plans for positioning, we propose a floor plan-assisted fusion positioning algorithm (FP-BP) using Bluetooth low energy (BLE) and pedestrian dead reckoning (PDR). In the considered system, a user holding a smartphone walks through a positioning area with BLE beacons installed on the ceiling, and can locate himself in real time. In particular, FP-BP consists of two phases. In the offline phase, FP-BP programmatically extracts map features from a stylized floor plan based on their binary masks, and constructs a mapping function to identify the corresponding map feature of any given position on the map. In the online phase, FP-BP continuously computes BLE positions and PDR results from BLE signals and smartphone sensors, where a novel grid-based maximum likelihood estimation (GML) algorithm is introduced to enhance BLE positioning. Then, a particle filter is used to fuse them and obtain an initial estimate. Finally, FP-BP performs post-position correction to obtain the final position based on its specific map feature. Experimental results show that FP-BP can achieve a real-time mean positioning accuracy of 1.19 m, representing an improvement of over 28% compared to existing floor plan-fused baseline algorithms.
Index Terms:
Bluetooth low energy (BLE), floor plan constraints, fusion positioning, pedestrian dead reckoning (PDR)
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
With the wide adoption of emerging applications such as smart cities, Industry 4.0, and extended reality (XR), the need for accurate indoor positioning has become more critical than ever. Although the global navigation satellite system (GNSS) has demonstrated high accuracy and wide coverage in outdoor scenarios, it still faces challenges in indoor applications, due to the weak satellite signals and environmental obstructions[2, 3]. To address these challenges, various indoor positioning technologies, such as WiFi[4], Bluetooth low energy (BLE)[5], radio frequency identification (RFID)[6], visible light positioning (VLP)[7], and ultra-wideband (UWB)[8], have been proposed. Among these, BLE positioning stands out due to its advantages of low cost and high energy efficiency[9].
@@
To improve the positioning accuracy of BLE, some fusion-based positioning technologies have been proposed. Among these, pedestrian dead reckoning (PDR) has emerged as a popular method[18]. Note that PDR typically comprises an inertial measurement unit (IMU) consisting of an accelerometer and a gyroscope, a magnetometer, and other sensors, and thus it can provide abundant extra information, such as movement speed and direction, over short periods. However, PDR-based positioning requires an accurate initial position and suffers from significant accuracy degradation over time due to cumulative errors of PDR[18, 19]. By integrating other positioning methods such as BLE, these errors can be effectively mitigated[13]. In particular, the absolute position references provided by BLE can help correct the PDR cumulative errors, and thus enhance the overall reliability of the fusion system.
@@ I-A Related Works
Current studies on fusing BLE/WiFi with PDR can be categorized into two types: (i) Calculate the BLE/WiFi positioning results first, typically through methods like trilateration[22],[23] or fingerprinting[20, 24], and then fuse these results with PDR; (ii) Fuse the received BLE/WiFi data, such as BLE RSSI or WiFi round-trip time (RTT), with IMU data directly[25]. For instance, Kong et al.[26] directly used an adaptive feedback extended Kalman filter (EKF) to fuse beacon positions, received RSSIs, and the result of PDR, to enhance the accuracy of the estimated position. In terms of filtering mechanisms, the above algorithms commonly employ the Kalman filter (KF)[22, 23] or its extended variations, such as the EKF[24, 26], as well as the particle filter (PF)[20]. However, none of the above traditional algorithms consider the integration of floor plans, which contain useful information such as environmental restrictions and pedestrian accessibility for positioning.
To further take the environmental factors into consideration, integrating floor plans in positioning systems has become a research hotspot[21, 27, 28, 29]. In general, the floor plan can also be regarded as a type of sensor[13]. In indoor positioning scenarios, floor plans are readily available[27] and can effectively constrain positioning results to the walkable area, thereby improving positioning performance[30, 31]. With respect to the filtering mechanisms, the PF is considered highly compatible with map information[21]. As such, many floor plan-integrated algorithms adopt PF for fusion. In[32], the authors generate particles in PF based on a specific angular probability density function (PDF), effectively avoiding “wall-crossing” particles. However, this algorithm requires frequent and extensive wall-distance checks, resulting in high computational costs in practical scenarios. The authors in[33] directly assign zero weights to “wall-crossing” particles in PF. Similarly, in[34], the authors use the map to remove particles that fall outside the map boundaries. However, these algorithms[33, 34] only simply eliminate particles in a position with large deviations, and this may lead to particle depletion and even positioning failure. The authors in[35] use changes in gyroscope yaws to determine whether the user has reached a corner, and thus can identify the corner and correct the yaw. However, this algorithm requires the pre-establishment of corner and fingerprint databases, which increases the offline workload. Although interesting, the above algorithms[33, 34, 35] have not fully utilized the floor plan information, such as the constraints of different obstacles, and therefore cannot provide stable positioning or require additional work.
@@
Although the above algorithms[20, 32, 33, 34, 35],[36, 37] and[38] have integrated floor plans, they struggle to fully utilize map information or adapt to complex indoor layouts, particularly when extracting map feature boundaries and leveraging their constraints in irregular environments. This may limit the feasibility and accuracy, or increase the deployment cost of the system in real-world scenarios. Thereby, a general and efficient floor plan-assisted fusion positioning algorithm deserves further investigation.
@@ I-B Contributions
The main contribution of this paper is a novel floor plan-assited real-time fusion positioning system. To the authors’ best knowledge, this is the first fusion framework that integrates a floor plan to achieve accurate real-time positioning without any limitations on map geometry. [^1] Based on the preliminary version[1] of this work, an enhanced BLE positioning algorithm named GML is proposed in this paper for the first time. Moreover, based on the proposed FP-BP algorithm, a prototype is developed for real-time positioning on mobile devices. Our key contributions are summarized as follows:
  • We propose a floor plan-assisted fusion positioning algorithm using BLE and PDR (FP-BP). When positioning, FP-BP fuses the BLE position and PDR through PF to obtain an initial result. Then, a post-position correction (PPC) algorithm is proposed to integrate the floor plan deeply into the system, using the map constructed from the floor plan in advance. It determines whether the initial result needs to be corrected by floor plan constraints, and applies corrections based on different map features to obtain the final position.
  • To programmatically build maps from various floor plans, we propose a feature extraction-based mapping method. This method first standardizes the floor plan style by filtering different obstacle areas and filling them with specific colors. Next, a binary mask-based method is applied to extract the inner contours of the floor plan. Finally, we establish coordinate systems on the floor plan and define a map feature function. This preprocessing mapping method can be completed in advance during the offline phase, and is widely applicable to various floor plans without any restriction.
  • To further enhance the accuracy of BLE positioning in FP-BP, we propose a grid-based maximum likelihood estimation (GML) algorithm. By analyzing the PDF of RSSI-estimated distance, we construct a likelihood function based on maximum likelihood estimation (MLE). Maximizing this likelihood function, the algorithm selects the optimal grid point within a virtual grid array as the estimated result, accommodating both dense and sparse beacon deployment scenarios. Finally, we apply mean filtering to multiple historical results to improve stability and accuracy.
Experimental results show that FP-BP algorithm can achieve a mean positioning accuracy of 1.19 m, which is improved by over 28% compared to existing floor plan-fused baselines. In addition, the proposed GML algorithm also reaches a mean accuracy of 2.01 m when the user keeps moving, which outperforms existing baselines by over 21%.
@@ I-C Notations and Paper Organization
We adopt the following notations: Matrices are denoted by uppercase boldface such as $\bm{R}$; Points and column vectors are denoted by lowercase boldface such as $\bm{v}$; $\bm{v}^{\mathsf{T}}$ and $\|\bm{v}\|$ represent the transposition and the norm of vector $\bm{v}$, respectively; Sets are denoted by calligraphic characters such as $\mathcal{A}$; $\{a_{i}\}_{i=1}^{n}$ represents a sequence containing $n$ elements and is arranged in order; $\mathcal{N}(\mu,\sigma^{2})$ represents a Gaussian distribution with an expectation of $\mu$ and a variance of $\sigma^{2}$; $p(\bm{x}|\bm{y})$ represents the PDF of a random variable $\bm{x}$ given the condition $\bm{y}$; and $\mathbb{E}_{\bm{x}|\bm{y}}[\bm{z}]$ denotes the expectation of a random variable $\bm{z}=\bm{z}(\bm{x})$ under the PDF $p(\bm{x}|\bm{y})$.
@@ II-A System Model
This section considers a typical positioning scenario as illustrated in Fig. 1, where a pedestrian carrying a smartphone passes through an indoor area and receives signals from BLE beacons for positioning. The beacons broadcast BLE signals at a fixed frequency, and the smartphone, i.e. the receiver, continuously scans for these signals. The smartphone also listens to the built-in sensors to obtain the user motion data, which is then fused with the BLE estimated position and the preprocessed floor plan information, allowing the user to locate himself in real time.
@@ III GML Algorithm for BLE Positioning
The expectation of $\widehat{d}_{i}$ can be then calculated as:
@@
where $c$ is a constant independent of $\bm{y}$. According to the MLE principle, the optimal estimated position $\bm{y}^{*}$ is determined by maximizing the $\ln\ell(\bm{y})$, which is equivalent to:
$\bm{y}^{*}=\mathop{\arg\min}_{\bm{y}\in\mathcal{G}^{\prime}}\sum_{i=1}^{N}\frac{1}{2\eta^{2}_{i}}\left(\log\widehat{d}_{i}-\log\|\bm{y}-\bm{b}^{(i)}\|\right)^{2},$ (19)
where $\mathcal{G}^{\prime}$ is the set of candidate points. Generally, (19) is a nonlinear optimization problem that is challenging to solve. We construct a set of grid points distributed throughout the entire area (see Section IV-A4), denoted as $\mathcal{G}$, and transform (19) into a problem of selecting an optimal point within a discrete grid point set $\mathcal{G}^{\prime}\subset\mathcal{G}$. Let $\mathrm{int}\mathcal{C}_{N}$ be the interior area of the convex hull $\mathcal{C}_{N}=\mathcal{CH}(\mathcal{B}_{N})$[41] formed by all beacons in $\mathcal{B}_{N}$, and $d(\bm{y}_{1},\bm{y}_{2})$ denote the Manhattan distance between point $\bm{y}_{1}$ and $\bm{y}_{2}$. When beacons are densely deployed in the positioning area, we can assume that the user’s position is always within $\mathrm{int}\mathcal{C}_{N}$, then the candidate grid points should also be close to the previous position $\bm{x}^{(k-1)}$:
$\mathcal{G}^{\prime}_{1}=\{\bm{y}\mid\bm{y}\in(\mathcal{G}\cap\mathrm{int}\mathcal{C}_{N})\wedge d(\bm{y},\bm{x}^{(k-1)})<d_{0}\}.$ (20)
@@
If the beacons are deployed sparsely, i.e., the $\mathrm{int}\mathcal{CH}(\mathcal{B})$ cannot geometrically cover the entire positioning area, the candidate grid points should not be restricted inside $\mathrm{int}\mathcal{C}_{N}$. In this case, $\mathcal{G}^{\prime}$ is represented as:
$\mathcal{G}^{\prime}_{2}=\{\bm{y}\mid\bm{y}\in\mathcal{G}\wedge d(\bm{y},\bm{x}^{(k-1)})<d_{0}\}.$ (21)
In the experimental scenario of this paper, $\mathcal{G}^{\prime}_{1}$ is taken as the candidate grid points set. For simplicity, it is assumed that the $\eta_{i}$ of all beacons are equal. The BLE positioning result at current time $t$ is obtained by averaging $\bm{y}^{*}$ and the previous $(n-1)$ BLE positioning results:
@@
From the above analysis, it is evident that GML algorithm takes the RSSI noise into consideration, and uses the posterior PDF of $\widehat{d}_{i}$ to form a unique likelihood, thereby reducing errors. To fuse BLE results into our positioning system, we will detail the proposed FP-BP algorithm in the following section.

### v3

@@ Abstract
Floor plans can provide valuable prior information that helps enhance the accuracy of indoor positioning systems. However, existing research typically faces challenges in efficiently leveraging floor plan information and applying it to complex indoor layouts. To fully exploit information from floor plans for positioning, we propose a floor plan-assisted fusion positioning algorithm (FP-BP) using Bluetooth low energy (BLE) and pedestrian dead reckoning (PDR). In the considered system, a user holding a smartphone walks through a positioning area with BLE beacons installed on the ceiling, and can locate himself in real time. In particular, FP-BP consists of two phases. In the offline phase, FP-BP programmatically extracts map features from a stylized floor plan based on their binary masks, and constructs a mapping function to identify the corresponding map feature of any given position on the map. In the online phase, FP-BP continuously computes BLE positions and PDR results from BLE signals and smartphone sensors, where a novel grid-based maximum likelihood estimation (GML) algorithm is introduced to enhance BLE positioning. Then, a particle filter is used to fuse them and obtain an initial estimate. Finally, FP-BP performs post-position correction to obtain the final position based on its specific map feature. Experimental results show that FP-BP can achieve a real-time mean positioning accuracy of 1.14 m, representing an improvement of over 29% compared to existing floor plan-fused baseline algorithms.
publicationid: pubid: 0000-$00$ © 2025 IEEE.
@@ I Introduction
With the wide adoption of emerging applications such as smart cities, Industry 4.0, and extended reality (XR), the need for accurate indoor positioning has become more critical than ever. Although the global navigation satellite systems (GNSSs) have demonstrated high accuracy and wide coverage in outdoor scenarios, they still face challenges indoors due to the weak satellite signals and environmental obstructions[2, 3]. To address these challenges, various indoor positioning technologies, such as WiFi[4], Bluetooth low energy (BLE)[5], radio frequency identification (RFID)[6], visible light positioning (VLP)[7], and ultra-wideband (UWB)[8], have been proposed. Among these, both WiFi and BLE share advantages such as low cost and wide availability in commercial devices. Nevertheless, BLE tends to offer additional practical advantages, owing to its lower power consumption[2, 9], higher deployment flexibility[9, 10] and generally improved positioning accuracy[11].
@@
To improve the positioning accuracy of BLE, some fusion-based positioning technologies have been proposed. Among these, pedestrian dead reckoning (PDR) has emerged as a popular method[18]. Note that PDR typically comprises an inertial measurement unit (IMU) consisting of an accelerometer and a gyroscope, a magnetometer, and other sensors, and thus it can provide abundant extra information, such as movement speed and direction, over short periods. However, PDR requires an accurate initial position and suffers from significant accuracy degradation over time due to cumulative errors of PDR[18, 19]. By integrating other positioning methods such as BLE, these errors can be effectively mitigated[10]. In particular, the absolute position references provided by BLE can help correct the PDR cumulative errors, and thus enhance the overall reliability of the fusion system.
@@ I-A Related Works
Current studies on fusing BLE/WiFi with PDR can be categorized into two types: (i) Calculate the BLE/WiFi positioning results first, typically through methods like trilateration[22, 23] or fingerprinting[20, 24], and then fuse these results with PDR; (ii) Fuse the received BLE/WiFi data, such as BLE RSSI or WiFi round-trip time (RTT), with IMU data directly[25]. For instance, Kong et al.[26] directly used an adaptive feedback extended Kalman filter (EKF) to fuse beacon positions, received RSSIs, and the result of PDR, to enhance the accuracy of the estimated position. In terms of filtering mechanisms, the above algorithms commonly employ the Kalman filter (KF)[22, 23] or its extended variations, such as the EKF[24, 26], as well as the particle filter (PF)[20]. However, none of the above traditional algorithms consider the integration of floor plans, which contain useful information such as environmental restrictions and pedestrian accessibility for positioning.
To further take the environmental factors into consideration, integrating floor plans in positioning systems has become a research hotspot[21, 27, 28, 29]. In general, the floor plan can also be regarded as a type of sensor[10]. In indoor scenarios, floor plans are readily available[27] and can effectively constrain positioning results to the walkable area, thereby improving system performance[30, 31]. With respect to the filtering mechanisms, the PF is highly compatible with map information[21]. As such, many floor plan-integrated algorithms adopt PF for fusion. In[32], the authors generate particles in PF based on a specific angular probability density function (PDF), effectively avoiding “wall-crossing” particles. However, this algorithm requires frequent and extensive wall-distance checks, resulting in high computational cost in practical scenarios. The authors in[29] and[33] directly assign zero weights to “wall-crossing” particles in PF. Similarly, in[34], the authors use the map to remove particles that fall outside the map boundaries. However, these algorithms[29, 33, 34] neglect implicit but crucial information such as the orientations of walls and corners. Moreover, the simple elimination of particles may lead to particle depletion or even positioning failure. The authors in[35] determine and identify whether the user has reached a specific corner through gyroscope data to help correct the yaw. However, this algorithm does not preprocess the floor plan and therefore cannot automatically extract the positions and orientations of corners. As a result, it requires the pre-establishment of a corner database, which inevitably increases the offline workload. Although interesting, the above algorithms[29, 33, 34, 35] have not fully utilized the floor plan information, and therefore cannot provide stable long-range positioning or require additional work.
@@
Although the above algorithms[20, 29, 32, 33, 34, 35, 36, 37] and[38] have integrated floor plans, they struggle to fully utilize map information or adapt to complex indoor layouts, particularly when extracting map feature boundaries and leveraging their constraints in irregular environments. This may limit the feasibility and accuracy, or increase the deployment cost of the system in real-world scenarios. Thereby, a general and efficient floor plan-assisted fusion positioning algorithm deserves further investigation.
@@ I-B Contributions
The main contribution of this paper is a novel floor plan-assisted fusion positioning system using BLE and PDR (FP-BP). To the authors’ best knowledge, this is the first PDR-based framework that integrates floor plans to achieve real-time positioning without limitations on map geometry. [^1] Based on the preliminary version[1] of this work, an enhanced BLE positioning algorithm named GML is proposed in this paper for the first time. Moreover, this paper extends the FP-BP algorithm to multi-floor scenarios to enhance its adaptability and feasibility. Furthermore, based on FP-BP, a prototype is developed for real-time positioning on mobile devices. Our key contributions are summarized as follows:
  • In the online phase, FP-BP first fuses BLE with PDR through the PF to obtain an initial estimation. Then, a post-position correction (PPC) mechanism is proposed to integrate the floor plan deeply into the system, using the map constructed from the floor plan in advance. This mechanism refines the initial estimation based on specific map features to obtain the final result, effectively exploiting the prior knowledge from the floor plan while preventing particle depletion.
  • In the offline phase, to programmatically build maps from various floor plans, we propose a feature extraction-based preprocessing method. The only manual step is to first standardize the floor plan by filling obstacle areas with distinguishable colors. Next, a binary mask-based method is applied to extract the inner contours of the floor plan. Finally, a map feature function is defined based on the pre-established planar coordinate systems. This method enables precise pixel-level recognition and extraction of map features, and is therefore widely applicable to various floor plans without restriction.
  • To further enhance BLE positioning accuracy, we propose a grid-based maximum likelihood estimation (GML) algorithm. By analyzing the posterior PDF of RSSI-estimated distances, we formulate an objective function based on maximum likelihood estimation (MLE) combined with a novel distance-based penalty term for refinement. Maximizing this function, the algorithm selects the optimal grid point within a virtual grid array as the estimated result, effectively accommodating both dense and sparse beacon deployment scenarios.
Experimental results show that FP-BP algorithm can achieve a mean positioning accuracy of 1.14 m, which is improved by over 29% compared to existing floor plan-fused baselines. In addition, the proposed GML algorithm also reaches a mean accuracy of 1.74 m when the user keeps moving, which outperforms existing non-fingerprinting baselines by over 34%.
@@ I-C Notations and Paper Organization
We adopt the following notations: Matrices are denoted by uppercase boldface such as $\bm{R}$; Points and column vectors are denoted by lowercase boldface such as $\bm{v}$; $\bm{v}^{\mathsf{T}}$ and $\|\bm{v}\|$ represent the transposition and the norm of vector $\bm{v}$, respectively; Sets are denoted by calligraphic characters such as $\mathcal{A}$; $\{a_{i}\}_{i=1}^{n}$ represents a sequence containing $n$ elements and is arranged in order; $\mathcal{N}(\mu,\sigma^{2})$ represents a Gaussian distribution with an expectation of $\mu$ and a variance of $\sigma^{2}$; $p(\bm{x}|\bm{y})$ represents the PDF of a random variable $\bm{x}$ given the condition $\bm{y}$; $\mathbb{E}_{\bm{x}|\bm{y}}[\bm{z}]$ denotes the expectation of a random variable $\bm{z}=\bm{z}(\bm{x})$ under the PDF $p(\bm{x}|\bm{y})$; $\mathbb{T}$ and $\mathbb{F}$ represent Boolean Truth and Falsehood values, respectively.
@@ II-A System Model
This section considers a typical positioning scenario as illustrated in Fig. 1, where a pedestrian carrying a smartphone passes through an indoor area and receives signals from BLE beacons for positioning. The beacons broadcast BLE signals at a fixed frequency, and the smartphone, i.e., the receiver, continuously scans for these signals. The smartphone also listens to the built-in sensors to obtain the user motion data, which is then fused with the BLE estimated position and the preprocessed floor plan information, allowing real-time active positioning for the user.
@@ III GML Algorithm for BLE Positioning
After obtaining the PDF $p(\widehat{d}_{i}|\bm{y})$, the expectation of $\widehat{d}_{i}$ can be then calculated as:
@@
where $c$ is a constant independent of $\bm{y}$. We denote the basic MLE function as:
$F_{\text{MLE}}(\bm{y})=\sum_{i=1}^{N}\frac{1}{\eta^{2}_{i}}\left(\log\widehat{d}_{i}-\log\|\bm{y}-\bm{b}^{(i)}\|\right)^{2}.$ (19)
According to the MLE principle[41], the optimal estimation is obtained by minimizing $F_{\text{MLE}}(\bm{y})$. However, such methods suffer from accuracy degradation due to the height difference between beacons and handheld devices. To mitigate this error, we further introduce a penalty term on the basis of (19) to enhance the influence of beacons with stronger RSSI on the estimation. We minimize the following objective function:
$\bm{y}^{*}=\mathop{\arg\min}_{\bm{y}\in\mathcal{G}^{\prime}}\left[F_{\text{MLE}}(\bm{y})+\kappa\sum_{i=1}^{N}\rho_{i}\|\bm{y}-\bm{b}^{(i)}\|\right],$ (20)
where $\kappa>0$ and $0<\rho_{i}<1$ are weighting parameters, and $\rho_{i}$ is defined to increase with the received RSSI $\widehat{R}_{i}$, which can be computed as follows:
$\rho_{i}=\mathop{\mathrm{softmax}}_{1\leq i\leq N}\tau\widehat{R}_{i},$ (21)
where $\tau>0$ is a tunable parameter. Moreover, in (20), $\mathcal{G}^{\prime}$ is the set of candidate points. Generally, (20) is a nonlinear optimization problem that is challenging to solve. We construct a set of grid points distributed throughout the entire area (see Section IV-A4), denoted as $\mathcal{G}$, and transform (20) into a problem of selecting an optimal point within a discrete grid point set $\mathcal{G}^{\prime}\subset\mathcal{G}$. Let $\mathrm{int}\mathcal{C}_{N}$ be the interior area of the convex hull $\mathcal{C}_{N}=\mathcal{CH}(\mathcal{B}_{N})$[42] formed by all beacons in $\mathcal{B}_{N}$, and $d(\bm{y}_{1},\bm{y}_{2})$ denote the Manhattan distance between point $\bm{y}_{1}$ and $\bm{y}_{2}$. When beacons are densely deployed in the positioning area, we can assume that the user’s position is always within $\mathrm{int}\mathcal{C}_{N}$, then the candidate grid points should also be close to the previous position $\bm{x}^{(k-1)}$:
$\mathcal{G}^{\prime}_{D}=\{\bm{y}\mid\bm{y}\in(\mathcal{G}\cap\mathrm{int}\mathcal{C}_{N})\wedge d(\bm{y},\bm{x}^{(k-1)})<d_{0}\}.$ (22)
@@
If the beacons are deployed sparsely, i.e., $\mathrm{int}\mathcal{CH}(\mathcal{B})$ cannot geometrically cover the entire area, $\mathrm{int}\mathcal{C}_{N}$ may fail to cover the actual user position, and thus may instead lead to positioning deviation. Therefore, we remove the convex hull constraint $\mathrm{int}\mathcal{C}_{N}$ and revise (22) accordingly:
$\mathcal{G}^{\prime}_{S}=\{\bm{y}\mid\bm{y}\in\mathcal{G}\wedge d(\bm{y},\bm{x}^{(k-1)})<d_{0}\}.$ (23)
In the experimental scenario of this paper, for simplicity, it is assumed that the $\eta_{i}$ of all beacons are equal[41] and normalized as 1. The BLE positioning result at current time $t$ is obtained by averaging $\bm{y}^{*}$ and the previous (n1)(n-1) BLE positioning results:
@@
From the above analysis, it is evident that GML algorithm takes the RSSI noise into consideration, uses the posterior PDF of $\widehat{d}_{i}$ to form a unique likelihood, and introduces a novel penalty term, thereby reducing errors. To fuse BLE results into our positioning system, we will detail the proposed FP-BP algorithm in the following section.